OpenStudy (anonymous):

28 is 80% of what number? how do you write this to figure it out

OpenStudy (anonymous):

28/x = 80/100

OpenStudy (anonymous):

do it using algebra let x = the number then 0.80 x = 28 divide both sides by 0.8 x = 28 / 0.8 = 35

OpenStudy (anonymous):

well if u have 80 out of 100 thats is 80% and u have 28/x that means it equals 35

OpenStudy (anonymous):

also u need to divide 4got to add that in there

myininaya (myininaya):

28=80%x 28=80x/100 28=20x/25 28=4x/5 28*5/4=x 7*5=x 35=x